OpenStudy (anonymous):

Solve the following system. y = x + 3 4x + y = 18 (6, 3) (3, 6) (-3, 6) (3, -6)

OpenStudy (johnweldon1993):

y = x + 3 4x + y = 18 Our first equation tells us what 'y' equals...so we can just plug that into the second equation for 'y' 4x + y = 18 becomes 4x + x + 3 = 18 Now simplify 4x + x + 3 = 18 -3 -3 ----------------- 4x + x = 15 5x = 15 -------- 5 5 x = 3 So if your 'x' = 3...what does 'y' equal?

OpenStudy (anonymous):

6 ?

OpenStudy (johnweldon1993):

Correct :)